Vector A has length 5 units and directed to the north. Vector b has length 8 units and is directed to the south. Calculate the magnitude and direction of a - b


1
Expert's answer
2020-09-14T10:19:53-0400
"\bold{b-a}=8-5=3"

The magnitude is 3.

The direction is to the south.